OpenStudy (anonymous):

Determine whether BC//DE. AD=15, DB=12, AE=10, and EC=8.

OpenStudy (anonymous):



OpenStudy (kc_kennylau):

Check if \(\dfrac{AD}{DB}=\dfrac{AE}{EC}\)

OpenStudy (anonymous):

ok hold up

OpenStudy (anonymous):

shouldn't it be AD/AE=DB/EC

OpenStudy (kc_kennylau):

Well they're the same aren't they

OpenStudy (anonymous):

well yeah but i got 120=120

ganeshie8 (ganeshie8):

By converse of side splitter theorem, if the line cuts two sides of a triangle proportionally, then it is parallel to the other side

ganeshie8 (ganeshie8):

since you got AD/DB = AE/EC, DE must be parallel to BC

OpenStudy (anonymous):

so then i did get it right

ganeshie8 (ganeshie8):

yup ! good job !!